# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ 2.4 Reducing $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
In the last part of the Technical Overview we will show that two problems that are believed to be hard for quantum computers, are solvable by a quantum computer with computable order interference, by reducing to the $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$ problem. In the first part of this section we show the reduction from $\mathsf{GI}$ which is more immediate, and in the second part we show the more involved reduction from $\mathsf{GapCVP}_{O\left(n^{2}\right)}$. To this end, we will first review the existing reduction from each of the problems to the more general $\mathsf{SD}_{\mathsf{poly}}$ problem (which we don’t know how to solve), and then move to our new reductions, to the easier problem of $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$.
@@
An input to the $\mathsf{GI}$ problem is a pair of (simple, undirected) graphs $\left(G_{0},G_{1}\right)$, and the problem is to decide whether the pair of graphs are isomorphic or not. More formally, for $n\in\mathbb{N}$ we define a graph as $G=\left([n],E\right)$, where $E\subseteq[n]\times[n]$ is the set of edges. Two graphs are isomorphic iff there exists an isomorphism between them, that is, a bijection (or, permutation) $f$ on $[n]$ such that for every pair $u,v\in[n]$: $\{u,v\}\in E_{0}\iff\{f(u),f(v)\}\in E_{1}$, where for $b\in\{0,1\}$, $E_{b}$ is the set of edges in the graph $G_{b}$.
As stated earlier in Section 2.1, there is an existing reduction from $\mathsf{GI}$ to the standard statistical difference problem $\mathsf{SD}_{\mathsf{poly}}$. More precisely, [GMW91] shows a statistical zero-knowledge protocol for the $\mathsf{GI}$ problem, out of which there can be derived the reduction $\mathsf{GI}\leq_{p}\mathsf{SD}_{0,1}$. The reduction to $\mathsf{SD}_{0,1}$ means that for an input pair $\left(G_{0},G_{1}\right)$ to the $\mathsf{GI}$ problem, if the graphs are isomorphic then the output pair of circuits $\left(C_{0},C_{1}\right)$ are such that the statistical distance between their output distributions is 00, and if the graphs are not isomorphic, the distance is $1$, which means that the supports of the output distributions are disjoint.
The reduction is as follows. Given the input graphs $\left(G_{0},G_{1}\right)$, the first circuit $C_{0}$ is defined such that for input randomness $r$ of polynomial length, it samples a uniformly random permutation $\sigma\in S_{n}$ on the elements $[n]$, applies $\sigma$ to the graph $G_{0}$, and the output of $C_{0}$ is the $\sigma$-permuted graph $\sigma\left(G_{0}\right)$ [^7] Recall that permuting $G$ means applying $\sigma$ to the vertices of $G$, but keeping the edges connecting the same vertex numbers.. The circuit $C_{1}$ is identical, only that it applies $\sigma$ to the second graph, $G_{1}$. To see the correctness of the reduction, recall that for two isomorphic graphs, their sets of all isomorphic graphs are equal, thus in that case, the circuits $C_{0},C_{1}$ sample from the exact same distribution (uniform distribution over their isomorphic graphs!) and the statistical distance is 00. In the other case where the graphs are not isomorphic, there does not exist a permutation $\sigma$ that will map one graph to the other, and in particular, the sets of permutations of the two graphs will never intersect, which will result in a statistical distance of $1$.
The reduction $\mathsf{GI}\leq_{p}\mathsf{SISD}_{\left(0,1,\log(n)\right)}$
@@
Our reduction follows by two arguments. The first step is that constant permutation mappings are efficiently computable and invertible. Formally, assume that we use adjacency matrix representation of a graph, so $n^{2}$ bits are sufficient to represent any $n$-vertex graph, then for a fixed permutation $\sigma\in S_{n}$, the circuit $C_{\sigma}:\{0,1\}^{n^{2}}\rightarrow\{0,1\}^{n^{2}}$ that permutes a given input graph $G$ with the permutation $\sigma$, is both efficiently computable and efficiently invertible. The second step is that a random permutation over $[n]$, which has $n!$ options, can be sampled in many small steps, each having a small number of options. Specifically, the Fisher-Yates shuffle [FY53], shows how to sample a random permutation in $n-1$ steps, each step $i=n,n-1,\cdots,3,2$ having $i$ options, and importantly, each step can be thought of as a fixed permutation by itself, which means the circuit that implements it is efficiently computable and invertible. Formally, for a decreasing $i=n,n-1,\cdots,3,2$, pick a random number $j\in[i]$, and swap between $i$ and $j$.
It remains to compile our observations into a $\log(n)$-invertible circuit sequence. Consider $n-1$ circuit pairs $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in\{n,n-1,\cdots,3,2\}}$, such that $C_{i,\rightarrow}:\{0,1\}^{n^{2}}\times\{0,1\}^{\lceil\log(n)\rceil}\rightarrow\{0,1\}^{n^{2}}$ maps between $n$-vertex graphs and uses the bits of randomness to pick a random element in $j\in[i]$, and then applies a swap between vertices $i$ and $j$. Per fixed randomness $j\in[i]$, the circuit $C_{i,\rightarrow}\left(\cdot,j\right)$ performs a swap between vertices $i,j$, which one can observe to be a fixed permutation.
There are some technicalities to deal with (like that sometimes ranges of randomness are represented by binary strings but are not of a size which is a power of $2$, which we use rejection sampling to deal with) in the full proof, but overall, the above circuit sequence implements the Fisher Yates shuffle. By generating one circuit sequence $\left(C^{0}_{i,\rightarrow},C^{0}_{i,\leftarrow}\right)_{i\in\{n,n-1,\cdots,3,2\}}$ to permute the graph $G_{0}$ and another $\left(C^{1}_{i,\rightarrow},C^{1}_{i,\leftarrow}\right)_{i\in\{n,n-1,\cdots,3,2\}}$ to permute $G_{1}$, both according to the Fisher Yates algorithm, which we break into small swap circuits, the correctness of the reduction follows similar lines to that of [GMW91]. The full proof is given in Section 6.
@@
The instance comes with a promise: It is either the case that the vector $\mathbf{t}$ is $d$-close to the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, or the vector is at least $\left(d\cdot g(n)\right)$-far from the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$. In simple words, a solving algorithm for $\mathsf{GapCVP}_{g(n)}$ gets the promise that the vector $\mathbf{t}$ is either close or far away from the lattice $\mathcal{L}_{\mathbf{B}}$, and needs to decide which is the case. The bigger the gap $g(n)$ is, the easier the problem becomes [^8] In particular, it is known that for an exponential $g(n)\approx 2^{n}$ the $\mathsf{GapCVP}_{2^{n}}$ problem is solvable in classical polynomial time.. What we will show is that for $g(n)=O\left(n^{2}\right)$ there is a reduction $\mathsf{GapCVP}_{g(n)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$. The inspiration and starting point of our reduction is the reduction $\mathsf{GapCVP}_{\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}}\leq_{p}\mathsf{SD}_{\mathsf{poly}}$ by Goldreich and Goldwasser [GG98], which we recall next.
@@
The [GG98] reduction shows how, for a smaller gap $\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$ (that is, a harder version of $\mathsf{GapCVP}$), we can reduce the Gap Closest Vector Problem to the Statistical Difference Problem $\mathsf{SD}$. Given a $\mathsf{GapCVP}_{g(n)}$ instance $\left(\mathbf{B},\mathbf{t},d\right)$ (for $g(n)=\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$), the reduction executes in classical polynomial time and outputs $\left(C_{0},C_{1}\right)$, such that the mapping should preserve $\prod_{\text{YES}}$ and $\prod_{\text{NO}}$ possession. Given the input, the [GG98] reduction first defines an upper bound $M:=M_{\left(\mathbf{B},\mathbf{t}\right)}:=2^{n}\cdot\max_{\mathbf{v}\in\{\mathbf{b}_{1},\cdots,\mathbf{b}_{n},\mathbf{t}\}}\left\lVert\mathbf{v}\right\rVert$. The upper bound is used to treat an inherently infinite lattice in a finite way, by considering only $M$-bounded coordinates. The definition of the circuits $C_{0}$, $C_{1}$ is as follows. The circuit $C_{0}$ first samples a uniformly random $M$-bounded coordinates vector $\mathbf{s}\in\mathbb{Z}_{M}^{n}$ and transforms the coordinates to a lattice vector $\mathbf{v}:=\mathbf{B}\cdot\mathbf{s}$. The final output of $C_{0}$ is a uniformly random point $\mathbf{p}$ inside the $n$-dimensional ball of radius $\frac{d\cdot g(n)}{2}$ around $\mathbf{v}$. The circuit $C_{1}$ is the same, with one change: At the end, we add $\mathbf{t}$ to $\mathbf{p}$. Note that if the distance of the vector $\mathbf{t}$ from the lattice is $d$, then it can be written as $\mathbf{t}=\mathbf{a}+\mathbf{e}$, such that $\mathbf{a}\in\mathcal{L}_{\mathbf{B}}$ and $\left\lVert\mathbf{e}\right\rVert=d$, such that $\mathbf{e}$ is the shortest vector which satisfies this equality.
@@
The rationale behind the reduction is that when having a uniform distribution over a lattice, then adding a lattice vector to the output keeps the distribution identical. This logic stays consistent for a distribution that samples balls around the lattice points. So, when looking at the change between $C_{0}$, $C_{1}$, the lattice component $\mathbf{a}$ in the vector $\mathbf{t}$ gets ”swallowed”, and the only vector that skews the distribution is $\mathbf{e}$. One can imagine how adding $\mathbf{e}$ pushes all balls around the lattice points in its direction and length, and this is exactly the move from the output distribution of $C_{0}$ to the output distribution of $C_{1}$. If $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then the skewing vector $\mathbf{e}$ has short length $\leq d$ compared to the radius $\frac{d\cdot g(n)}{2}$ of the balls, which makes the balls have significant intersection, which in turn implies an upper bounded statistical distance between the two circuits (specifically, the statistical distance is $\leq 1-\frac{1}{n^{c}}$ for some constant $c>0$) and thus $\left(C_{0},C_{1}\right)\in\prod_{\text{YES}}$. If $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$ and $\mathbf{e}$ has length $>d\cdot g(n)$, one can observe how it pushes the $n$-balls (which have radius $\frac{d\cdot g(n)}{2}$, which is in turn half of the distance to the vector $\mathbf{e}$) to be completely disjoint from the original ones, which implies a statistical distance of $1$ and thus $\left(C_{0},C_{1}\right)\in\prod_{\text{NO}}$.
@@
The [GG98] reduction maps to circuits $\left(C_{0},C_{1}\right)$ that sample a uniform point $\mathbf{p}$ inside the radius-$\frac{d\cdot g(n)}{2}$, $n$-dimensional ball, around a uniform lattice vector $\mathbf{v}$. While sampling a random lattice vector can be done in a sequentially-invertible manner, sampling a uniform point inside the $n$-ball by a sequentially invertible algorithm seems like a fundamental problem. More specifically (but still in a nutshell), the reason that sampling from the $n$-ball seems hard, is that known algorithmic techniques to get a uniformly random sample from the $n$-ball include normalization. That is, we sample from some large set $S$, vectors that are not necessarily inside the wanted radius of the $n$-ball, and then normalize. The problem arises because normalization is a function of the entire vector, and it is inherently a non-invertible function (for example, all vectors of the same direction but different lengths, will map to the same normalized vector).
@@
  • The distribution $P$ should not have a preferred direction, out of the $n$ directions in $\mathbb{Z}^{n}$. This requirement is to deal with arbitrary locations of the vector $\mathbf{t}$ with respect to the lattice $\mathcal{L}_{\mathbf{B}}$ (that is, if $P$ that the reduction picks prefers some direction $i\in[n]$, then the vector $\mathbf{t}$ can be chosen as a function of that direction, tricking the reduction).
@@
  • We imagine the support of $P$ as an $n$-dimensional shape. The distribution $P$ should be as dense as possible in its shape. The density requirement emerges from the case where $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ – in that case we sample uniform samples from the shape around the lattice vectors, and the more the shape is sparse, the less our ability to give a good upper bound on the total variation distance between the output distributions of the circuits $C_{0}$, $C_{1}$.
@@
It turns out that the distribution which maximizes the above two requirements is indeed the uniform distribution inside the $n$-ball. The challenge in our reduction is to find a proper distribution $P$ to sample from, and then design a $O\left(\log(n)\right)$-sequentially-invertible sampling algorithm. In fact, we will show a distribution $P$ with a $1$-sequentially-invertible sampling algorithm.
The distribution $P$.
@@
The distribution $P$ we choose is the $n$-dimensional discrete truncated Gaussian distribution $D^{n}_{B}$ (as per Definition 3.12, $B\in\mathbb{N}$ is the truncation parameter, assume it to be $\approx\frac{d\cdot g(n)}{2\cdot\sqrt{n}}$). It is a known fact that $D^{n}_{B}$ is invariant under rotations in $n$-space, which exactly means it does not prefer any direction. In terms of sparsity, the Gaussian distribution is obviously less sparse than the uniform $n$-ball, which is (part of) the reason that we need a bigger gap $g(n)$ for which we know how to reduce $\mathsf{GapCVP}_{g(n)}$ to $\mathsf{SISD}$, compared to the better gap $\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$ from the previous reduction. As part of the full proof of the reduction (given in Section 7) we analyze the Gaussian distribution’s sparsity in this manner, and the loss of hardness (which translates to the larger $g(n)\approx n^{2}>>\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$) that the sparsity incurs. It remains to show a sequentially-invertible sampling algorithm for $P:=D^{n}_{B}$. One of the basic properties of the Gaussian distribution says that sampling from $D^{n}_{B}$ can be done by sampling $n$ i.i.d. samples from $D_{B}$ (one sample for each of the coordinates). All we need to find is a sequentially-invertible algorithm to sample from the discrete Gaussian distribution $D_{B}$.
@@
Sampling from $D_{B}$ has a plethora of existing techniques and algorithms (e.g., [GPV08, Pei10, MP12, Ros19]), but it is unclear how to show that these can be executed by a sequentially invertible algorithm. In fact, we did not find any sequentially invertible algorithm that samples from a distribution with negligible total variation distance from $D_{B}$. We give the intuitive steps behind our solution in the following.

### v2

@@ 2.4 Reducing $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
In the last part of the Technical Overview we will show that two problems that are believed to be hard for quantum computers, are solvable by a quantum computer with computable order interference, by reducing to the $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$ problem. In the first part of this section we show the reduction from $\mathsf{GI}$ which is more immediate, and in the second part we show the more involved reduction from $\mathsf{GapCVP}_{O\left(n^{2}\right)}$. To this end, we will first review the existing reductions from each of the problems to the more general $\mathsf{SD}_{\mathsf{poly}}$ problem (which we don’t know how to solve), and then move to our new reductions, to the easier problem of $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$.
@@
An input to the $\mathsf{GI}$ problem is a pair of (simple, undirected) graphs $\left(G_{0},G_{1}\right)$, and the problem is to decide whether the pair of graphs is isomorphic or not. More formally, for $n\in\mathbb{N}$ we define a graph as $G=\left([n],E\right)$, where $E\subseteq[n]\times[n]$ is the set of edges. Two graphs are isomorphic iff there exists an isomorphism between them, that is, a bijection (or, permutation) $f$ on $[n]$ such that for every pair $u,v\in[n]$: $\{u,v\}\in E_{0}\iff\{f(u),f(v)\}\in E_{1}$, where for $b\in\{0,1\}$, $E_{b}$ is the set of edges in the graph $G_{b}$.
As stated earlier in Section 2.1, there is an existing reduction from $\mathsf{GI}$ to the standard statistical difference problem $\mathsf{SD}_{\mathsf{poly}}$. More precisely, [GMW91] shows a statistical zero-knowledge protocol for the $\mathsf{GI}$ problem, out of which there can be derived the reduction $\mathsf{GI}\leq_{p}\mathsf{SD}_{a,b}$, for $a=0$, $b=1$. The reduction to $\mathsf{SD}_{0,1}$ means that for an input pair $\left(G_{0},G_{1}\right)$ to the $\mathsf{GI}$ problem, if the graphs are isomorphic then the output pair of circuits $\left(C_{0},C_{1}\right)$ are such that the statistical distance (another term for total variation distance) between their output distributions is 00, and if the graphs are not isomorphic, the distance is $1$, which means that the supports of the output distributions are disjoint.
The reduction is as follows. Given the input graphs $\left(G_{0},G_{1}\right)$, the first circuit $C_{0}$ is defined such that for input randomness $r$ of polynomial length, it samples a uniformly random permutation $\sigma\in S_{n}$ on the elements $[n]$, applies $\sigma$ to the graph $G_{0}$, and the output of $C_{0}$ is the $\sigma$-permuted graph $\sigma\left(G_{0}\right)$ [^8] Recall that permuting $G$ means applying $\sigma$ to the vertices of $G$, but keeping the edges connecting the same vertex numbers.. The circuit $C_{1}$ is identical, only that it applies $\sigma$ to the second graph, $G_{1}$. To see the correctness of the reduction, recall that for two isomorphic graphs, their sets of all isomorphic graphs are equal, thus in that case, the circuits $C_{0},C_{1}$ sample from the exact same distribution (uniform distribution over their isomorphic graphs!) and the statistical distance is 00. In the other case where the graphs are not isomorphic, there does not exist a permutation $\sigma$ that will map one graph to the other (because such permutation will act as an isomorphism between the graphs), and in particular, the sets of permutations of the two graphs will never intersect, which will result in a statistical distance of $1$.
Our reduction $\mathsf{GI}\leq_{p}\mathsf{SISD}_{\left(0,1,\log(n)\right)}$
@@
Our reduction follows by two arguments. The first step is that constant permutation mappings are efficiently computable and invertible. Formally, assume that we use adjacency matrix representation of a graph, so $n^{2}$ bits are sufficient to represent any $n$-vertex graph. Then, for a fixed permutation $\sigma\in S_{n}$, the circuit $C_{\sigma}:\{0,1\}^{n^{2}}\rightarrow\{0,1\}^{n^{2}}$ that permutes a given input graph $G$ with the permutation $\sigma$, is both efficiently computable and efficiently invertible. The second step is that a random permutation over $[n]$, which has $n!$ options, can be sampled in many small steps, each having a small (that is, polynomially-bounded) number of options, and crucially, each step is invertible. Specifically, the Fisher-Yates shuffle [FY53], shows how to sample a random permutation in $n-1$ steps, each step $i=n,n-1,\cdots,3,2$ having $i$ options, and importantly, each step is a fixed permutation in of itself, which means the circuit that implements it is efficiently computable and invertible. Formally, for a decreasing $i=n,n-1,\cdots,3,2$, pick a random number $j\in[i]$, and swap between $i$ and $j$.
It remains to compile our observations into a $\log(n)$-invertible circuit sequence. Consider $n-1$ circuit pairs $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in\{n,n-1,\cdots,3,2\}}$, such that $C_{i,\rightarrow}:\{0,1\}^{n^{2}}\times\{0,1\}^{\lceil\log(n)\rceil}\rightarrow\{0,1\}^{n^{2}}$ maps between $n$-vertex graphs and uses the bits of randomness to pick a random element in $j\in[i]$, and then applies a swap between vertices $i$ and $j$. Per fixed randomness $j\in[i]$, the circuit $C_{i,\rightarrow}\left(\cdot,j\right)$ performs a swap between vertices $i,j$, which one can observe to be a fixed permutation. By generating a circuit sequence $\left(C^{b}_{i,\rightarrow},C^{b}_{i,\leftarrow}\right)_{i\in\{n,n-1,\cdots,3,2\}}$ to permute the graph $G_{b}$ (for $b\in\{0,1\}$), where both sequences are generated as above according to the Fisher-Yates algorithm, the rest of the correctness of the reduction then follows similar lines to that of [GMW91]. The full proof is given in Section 6.
@@
The input comes with a promise: It is either the case that the vector $\mathbf{t}$ is $d$-close to the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, or the vector is farther than $d\cdot g(n)$ from the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$. In simple words, a solving algorithm for $\mathsf{GapCVP}_{g(n)}$ gets the promise that the vector $\mathbf{t}$ is either close or far away from the lattice $\mathcal{L}_{\mathbf{B}}$, and needs to decide which is the case. The bigger the gap $g(n)$ is, the easier the problem becomes [^9] In particular, it is known that for an exponential $g(n)\approx 2^{n}$ the $\mathsf{GapCVP}_{2^{n}}$ problem is solvable in classical polynomial time [AKS01].. What we will show is that for $g(n)=O\left(n^{2}\right)$ there is a reduction $\mathsf{GapCVP}_{g(n)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$. The inspiration and starting point of our reduction is the reduction $\mathsf{GapCVP}_{\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}}\leq_{p}\mathsf{SD}_{\mathsf{poly}}$ by Goldreich and Goldwasser [GG98], which we recall next.
@@
The [GG98] reduction shows how, for a gap $\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$, we can reduce the Gap Closest Vector Problem to the Statistical Difference Problem $\mathsf{SD}$. Given a $\mathsf{GapCVP}_{g(n)}$ instance $\left(\mathbf{B},\mathbf{t},d\right)$ (for $g(n)=\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$), the reduction executes in classical polynomial time and outputs $\left(C_{0},C_{1}\right)$, such that in case the $\mathsf{GapCVP}$ input is in $\prod_{\text{YES}}$ then the statistical distance between the output distributions of the circuits is bounded by $1-\frac{1}{n^{c}}$ (for some constant $c\in\mathbb{N}$), and in case the $\mathsf{GapCVP}$ input is in $\prod_{\text{NO}}$ then the statistical distance between the output distributions of the circuits is the maximal $1$. Given the input, the [GG98] reduction first defines an upper bound $M:=M_{\left(\mathbf{B},\mathbf{t}\right)}:=2^{n}\cdot\max_{\mathbf{v}\in\{\mathbf{b}_{1},\cdots,\mathbf{b}_{n},\mathbf{t}\}}\left\lVert\mathbf{v}\right\rVert$. The upper bound is used to treat an inherently infinite lattice in a finite way, by considering only $M$-bounded coordinates. The definition of the circuits $C_{0}$, $C_{1}$ is as follows. The circuit $C_{0}$ first samples a uniformly random $M$-bounded coordinates vector $\mathbf{s}\in\mathbb{Z}_{M}^{n}$ and transforms the coordinates to a lattice vector $\mathbf{v}:=\mathbf{B}\cdot\mathbf{s}$. The final output of $C_{0}$ is a uniformly random point $\mathbf{p}$ inside the $n$-dimensional ball of radius $\frac{d\cdot g(n)}{2}$ around $\mathbf{v}$. The circuit $C_{1}$ is the same, with one change: At the end, we add $\mathbf{t}$ to $\mathbf{p}$. Note that if the distance of the vector $\mathbf{t}$ from the lattice is $d$, then it can be written as $\mathbf{t}=\mathbf{a}+\mathbf{e}$, such that $\mathbf{a}\in\mathcal{L}_{\mathbf{B}}$ and $\left\lVert\mathbf{e}\right\rVert=d$, such that $\mathbf{e}$ is the shortest vector which satisfies this equality.
@@
The rationale behind the reduction is that when having a uniform distribution over a lattice, then adding a lattice vector to the output keeps the distribution identical. This logic stays consistent for a distribution that samples balls around the lattice points. So, when looking at the change between $C_{0}$, $C_{1}$, the lattice component $\mathbf{a}$ in the vector $\mathbf{t}$ gets ”swallowed”, and the only vector that skews the distribution is $\mathbf{e}$. One can imagine how adding $\mathbf{e}$ pushes all $n$-balls around the lattice points in its direction and length, and this is exactly the move from the output distribution of $C_{0}$ to the output distribution of $C_{1}$. If $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then the skewing vector $\mathbf{e}$ has short length $\leq d$ compared to the radius $\frac{d\cdot g(n)}{2}$ of the balls, which makes the balls have significant intersection, which in turn implies an upper bounded statistical distance between the output distributions of the two circuits (specifically, it can be proven that the statistical distance is $\leq 1-\frac{1}{n^{c}}$ for some constant $c\in\mathbb{N}$) and thus $\left(C_{0},C_{1}\right)\in\prod_{\text{YES}}$. If $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$ and $\mathbf{e}$ has length $>d\cdot g(n)$, one can observe how it pushes the $n$-balls (which have radius $\frac{d\cdot g(n)}{2}$, which is in turn half of the distance to the vector $\mathbf{e}$) to be completely disjoint from the original ones, which implies a statistical distance of $1$ and thus $\left(C_{0},C_{1}\right)\in\prod_{\text{NO}}$.
@@
The [GG98] reduction maps to circuits $\left(C_{0},C_{1}\right)$ that sample a uniform point $\mathbf{p}$ inside the radius-$\frac{d\cdot g(n)}{2}$, $n$-dimensional ball, around a uniform lattice vector $\mathbf{v}$. While sampling a random lattice vector can be done in a sequentially-invertible manner, sampling a uniform point inside the $n$-ball by a sequentially invertible algorithm seems like a fundamental problem. More specifically (but still in a nutshell), the reason that sampling from the $n$-ball seems hard, is that known algorithmic techniques to get a uniformly random sample from the $n$-ball rely on normalization. That is, we sample from some large set $S$, vectors that are not necessarily inside the wanted radius of the $n$-ball, and then normalize. The problem arises because normalization is a function of the entire vector, and it is inherently a non-invertible function (for example, all vectors of the same direction but different lengths, will map to the same normalized vector).
@@
  • $n$-dimensional symmetry: The distribution $P$ should not have a preferred direction, out of the $n$ directions in $\mathbb{Z}^{n}$. This requirement is to deal with arbitrary locations of the vector $\mathbf{t}$ with respect to the lattice $\mathcal{L}_{\mathbf{B}}$ (that is, if the $P$ we pick prefers some direction $i\in[n]$, then the vector $\mathbf{t}$ can be chosen as a function of that direction, tricking the reduction).
@@
  • Density: We imagine the support of $P$ as an $n$-dimensional shape, by looking at the elements that are farthest away from the origin and thinking on them as a boundary for the shape. The distribution $P$ should be as dense as possible in its shape. The density requirement emerges from the case where $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ – in that case we sample uniform samples from the shape around the lattice vectors, and the more the shape is sparse, the less our ability to give a good upper bound on the total variation distance between the output distributions of the circuits $C_{0}$, $C_{1}$.
@@
It turns out that the distribution which maximizes the above two requirements is indeed the uniform distribution inside the $n$-ball. The challenge in our reduction is to find a proper distribution $P$ that satisfies the above, and also has a $O\left(\log(n)\right)$-sequentially-invertible sampling algorithm. In fact, we will show a distribution $P$ with a $1$-sequentially-invertible sampling algorithm.
A sequentially-invertible distribution $P$.
@@
The distribution $P$ we choose is the $n$-dimensional discrete truncated Gaussian distribution $D^{n}_{B}$ (as per Definition 3.12, $B\in\mathbb{N}$ is the truncation parameter, assume it to be $\approx\frac{d\cdot g(n)}{2\cdot\sqrt{n}}$). It is a known fact that $D^{n}_{B}$ is invariant under rotations in $n$-space, which exactly means it does not prefer any direction of the $n$ directions in $\mathbb{Z}^{n}$. In terms of density, the Gaussian distribution is obviously less dense in its shape than the uniform $n$-ball, which is (part of) the reason that we need a bigger gap $g(n)=O\left(n^{2}\right)$ for which we know how to reduce $\mathsf{GapCVP}_{g(n)}$ to $\mathsf{SISD}$, compared to the better gap $\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$ from the previous reduction. As part of the full proof of the reduction (given in Section 7) we analyze the Gaussian distribution’s sparsity in this manner, and the loss of hardness (which translates to the larger $g(n)\approx n^{2}>>\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$) that the sparsity incurs.
It remains to show a sequentially-invertible sampling algorithm for $P:=D^{n}_{B}$. One of the basic properties of the Gaussian distribution says that sampling from $D^{n}_{B}$ can be done by sampling $n$ i.i.d. samples from $D_{B}$ (one sample for each of the coordinates). All we need to find is a sequentially-invertible algorithm to sample from the discrete truncated Gaussian distribution $D_{B}$.
@@
Sampling from $D_{B}$ has a plethora of existing techniques and algorithms (e.g., [GPV08, Pei10, MP12, Ros19]), but it is unclear how to show that these can be executed by a sequentially invertible algorithm. In fact, we did not find any sequentially invertible algorithm that samples from a distribution with negligible total variation distance to $D_{B}$. Once we relax our needs to sample from the discrete Gaussian distribution only approximately (that is, with a non-negligible total variation distance), we find a solution. We give the intuitive steps behind our solution in the following.